# Explicit Runge-Kutta Methods with MQ and IMQ-Radial Basis Functions

### v1

@@ 3.2 Three-stage forth-order methods
$\tau_{n}=$ $\frac{u_{n+1}-u_{n}}{h}-(b_{1}k_{1}+b_{2}k_{2}+b_{3}k_{3})$
$=(1-b_{1}-b_{2}-b_{3})f+h\left(-a_{21}b_{2}ff_{u}-a_{31}b_{3}ff_{u}-a_{32}b_{3}ff_{u}-b_{2}c_{2}f_{t}-b_{3}c_{3}f_{t}+\frac{ff_{v}}{2}+\frac{f_{t}}{2}\right)$
$+h^{2}\biggl{(}-\frac{a_{21}^{2}b_{2}f^{2}f_{uu}}{2}-a_{21}a_{32}b_{3}ff_{u}^{2}-a_{21}b_{2}c_{2}ff_{ut}-a_{31}a_{32}b_{3}f^{2}f_{uu}-a_{31}b_{3}c_{3}ff_{ut}-\frac{a_{32}^{2}b_{3}f^{2}f_{uu}}{2}$
$-a_{32}b_{3}c_{2}f_{t}f_{v}-a_{32}b_{3}c_{2}f_{t}f_{u}-a_{32}b_{3}c_{3}ff_{ut}+\frac{\epsilon_{n2}^{2}b_{2}c_{2}^{2}u_{n}f_{u}}{2}-\frac{b_{2}c_{2}^{2}f_{tt}}{2}+\frac{\epsilon_{n3}^{2}b_{3}c_{3}^{2}u_{n}}{2}-\frac{b_{3}c_{3}^{2}f_{tt}}{2}$
$+\frac{f^{2}f_{uu}}{6}+\frac{ff_{u}^{2}}{6}+\frac{ff_{ut}}{3}+\frac{f_{t}f_{u}}{6}+\frac{f_{tt}}{6}\biggl{)}+h^{3}\biggl{(}-\frac{-a_{21}^{3}b_{2}f^{3}f_{uuu}}{6}-\frac{a_{21}^{2}a_{32}b_{3}f^{2}f_{u}f_{uu}}{2}$
$-a_{21}b_{2}c_{2}f^{2}f_{uut}-a_{21}a_{31}a_{32}b_{3}f^{2}f_{u}f_{uu}-a_{21}a_{32}^{2}b_{3}f^{2}f_{u}f_{uu}-a_{21}a_{32}b_{3}c_{2}ff_{u}f_{ut}$
$-a_{21}a_{32}b_{3}c_{3}ff_{u}f_{ut}+\frac{a_{21}b_{2}c_{2}^{2}u_{n}\epsilon_{n2}^{2}ff_{uu}}{2}-\frac{a_{21}b_{2}c_{2}^{2}\epsilon_{n2}^{2}ff_{u}}{2}-a_{21}b_{2}c_{2}^{2}ff_{utt}-\frac{a_{31}^{3}b_{3}f^{3}f_{uuu}}{6}$
$-\frac{a_{31}^{2}a_{32}b_{3}f^{3}f_{uuu}}{2}-\frac{a_{31}^{2}b_{3}c_{3}f^{2}f_{uut}}{2}-\frac{a_{31}a_{32}^{2}b_{3}f^{3}f_{uuu}}{2}-a_{31}a_{32}b_{3}c_{2}ff_{t}f_{uu}-\frac{a_{31}b_{3}c_{3}^{2}u_{n}\epsilon_{3n}^{2}ff_{uu}}{2}$
$-\frac{a_{31}b_{3}c_{3}^{2}\epsilon_{3n}^{2}ff_{u}}{2}-\frac{a_{31}b_{3}c_{3}^{2}ff_{utt}}{2}-a_{32}^{2}b_{3}c_{2}ff_{t}f_{uu}-\frac{a_{32}^{2}b_{3}c_{3}f^{2}f_{uut}}{2}+\frac{a_{32}b_{3}c_{2}^{2}u_{n}\epsilon_{2n}^{2}f_{u}^{2}}{2}-\frac{a_{32}b_{3}c_{2}^{2}f_{tt}f_{u}}{2}$
$-a_{32}b_{3}c_{2}c_{3}f_{t}f_{u}t-\frac{a_{32}b_{3}c_{3}^{3}u_{n}\epsilon_{3n}^{2}ff_{uu}}{2}-\frac{a_{32}b_{3}c_{3}^{2}ff_{utt}}{2}+\frac{b_{2}c_{2}^{2}u_{n}\epsilon_{2}^{2}f_{ut}}{2}-\frac{b_{2}c_{2}^{3}f}{6}+\frac{b_{3}c_{3}^{3}u_{n}f_{ut}\epsilon_{3n}^{2}}{2}$
$-\frac{b_{3}c_{3}^{3}f_{ttt}}{6}+\frac{f^{3}f_{uuu}}{24}+\frac{f^{2}f_{u}f_{u}u}{6}+\frac{f^{2}f_{vt}}{8}+\frac{ff_{t}f_{u}u}{8}+\frac{ff_{u}^{3}}{24}+\frac{5ff_{u}f_{ut}}{24}+\frac{ff_{ut}}{8}$
$+\frac{f_{t}f_{u}^{2}}{24}+\frac{f_{t}f_{ut}}{8}+\frac{f_{u}f_{tt}}{24}+\frac{f_{ttt}}{24}\biggl{)}+\mathcal{O}(h^{4}).$
@@
    $a_{21}=c_{2}=\frac{1}{2},a_{31}=-1,a_{32}=2,c_{3}=1,b_{1}=\frac{1}{6},b_{2}=\frac{2}{3},b_{3}=\frac{1}{6},$
    $\epsilon^{2}_{n2}=\frac{-f^{2}f_{u}f_{uu}-ff_{t}f_{uu}+ff_{u}^{3}-ff_{u}f_{tu}+f_{t}f_{u}^{2}-f_{t}f_{tu}}{uff_{uu}-uf_{u}^{2}+uf_{tu}-ff_{u}},$
    $\epsilon^{2}_{n3}=-\epsilon^{2}_{n2}.$
@@
    $a_{21}=c_{2}=\frac{1}{3},a_{31}=-\frac{5}{12},a_{32}=\frac{5}{4},c_{3}=\frac{5}{6},b_{1}=\frac{1}{10},b_{2}=\frac{1}{2},b_{3}=\frac{2}{5},$
    $\epsilon^{2}_{n2}$ $=\frac{-ff_{t}f_{uu}+3ff_{u}^{3}+ff_{u}f_{tu}+3f_{t}f_{u}^{2}-f_{t}f_{tu}+f_{tt}f_{u}}{uff_{uu}-2uf_{u}^{2}+uf_{tu}-ff_{u}},$
    $\epsilon^{2}_{n3}$ $=-\frac{1}{5}\epsilon^{2}_{n2}.$
@@
    $a_{21}=c_{2}=1,a_{31}=\frac{1}{4},a_{32}=\frac{1}{4},c_{3}=\frac{1}{2},b_{1}=\frac{1}{6},b_{2}=\frac{1}{6},b_{3}=\frac{2}{3},$
    $\epsilon^{2}_{n2}$ $=\frac{-ff_{t}f_{uu}-ff_{u}^{3}+ff_{u}f_{tu}-f_{t}f_{u}^{2}-f_{t}f_{tu}+f_{tt}f_{u}}{uff_{uu}+2vf_{u}^{2}+uf_{tu}-ff_{u}},$
    $\epsilon^{2}_{n3}$ $=-\epsilon^{2}_{n2}.$
@@
    $a_{21}=c_{2}=\frac{1}{2},a_{31}=0,a_{32}=c_{3}=\frac{3}{4},c_{3}=1,b_{1}=\frac{2}{9},b_{2}=\frac{1}{3},b_{3}=\frac{4}{9},$
    $\epsilon^{2}_{n2}$ $=\frac{\frac{1}{3}f^{3}f_{uuu}+f^{2}f_{tuu}+4ff_{u}^{3}+ff_{ttu}+4f_{t}f_{u}^{2}+\frac{1}{3}f_{ttt}}{uff_{uu}-4uf_{u}^{2}+uf_{tu}-ff_{u}},$
    $\epsilon^{2}_{n3}$ $=-\frac{1}{3}\epsilon^{2}_{n2}.$
4 Convergence of RBF RUNGE-KUTTA Method
Numerical methods derived in previous section can be applied on IVP if there exists unique solution. IVP of the form (1.1) has a unique solution if $f(t,u)$ satisfies Lipschitz continuity with respect to second-coordinate that is $|f(t,u_{1})-f(t,u_{2})|\leq L|u_{1}-u_{2}|$, where $L$ is Lipschitz constant. Now, we prove the convergence of above proposed methods. Note that, we solve the IVP on an interval $[a,b]$ which is divided by uniform mesh length $h=\dfrac{b-a}{N}$ as $N+1$ points with mesh grid points $a=t_{0}<t_{1}<...<t_{N}=b$ where $t_{n}=a+nh,n=0,1,...,N.$ We also denote $u_{n}=u(t_{n})$ as exact solution and $v_{n}$ as numerical approximation for the proposed RBF Runge-Kutta schemes.
@@ 4 Convergence of RBF RUNGE-KUTTA Method
Suppose, $E_{n+1}$ be the error between $u_{n+1}$ and $v_{n+1}$ for fix $t=t_{n}$. Now,
@@
$\leq L\biggl{|}\left(u_{n}+ha_{21}k_{1}(u_{n})\right)\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}-\left(v_{n}+ha_{21}k_{1}(v_{n})\right)\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}\biggl{|}$
$\leq L\biggl{|}(u_{n}-v_{n})\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}+a_{21}hL(u_{n}-v_{n})\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\biggl{|}$
@@
$L$ is Lipschitz constant.
$E_{n+1}$ $=|u_{n+1}-v_{n+1}|$
$=|(u_{n}-v_{n})+h[b_{1}(k_{1}(u_{n})-k_{1}(v_{n}))+b_{2}(k_{2}(u_{n})-k_{2}(v_{n}))]+h\tau_{n}|$
$\leq|u_{n}-v_{n}|+h[b_{1}|k_{1}(u_{n})-k_{1}(v_{n})|+b_{2}|k_{2}(u_{n})-k_{2}(v_{n})|]+h|\tau_{n}|$
$\leq\left(1+hLb_{1}+b_{2}(1+hL)\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\right)|u_{n}-v_{n}|+h|\tau_{n}|$
$=\phi_{n}E_{n}+h|\tau_{n}|$
@@
$\phi_{n}$ $=1+b_{1}hL+b_{2}(1+a_{21}hL)\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}$
$\leq 1+hL+e^{hL}\mathcal{C}_{2}h$
$\leq e^{hL}+e^{hL}\mathcal{C}_{2}h$
$=e^{hL}(1+\mathcal{C}_{2}h)$
$=e^{h(L+\mathcal{C}_{2})}$
@@
$\mathcal{C}_{2}=\sup_{\begin{subarray}{c}0<h\leq b-a\\ n=0,\dots,N-1\end{subarray}}\frac{\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}}{h},$
@@
$-f\left(t_{n},(v_{n}+h(a_{31}k_{1}(v_{n})+a_{32}k_{2}(v_{n}))\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\right)\bigg{|}$
$\leq L\left|(u_{n}-v_{n})+\left(ha_{31}L|u_{n}-v_{n}|+ha_{32}L(1+hL)\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}|u_{n}-v_{n}|\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\right)\right|$
@@
$L$ is Lipschitz constant.
$E_{n+1}$ $=|u_{n+1}-v_{n+1}|$
$=|(u_{n}-v_{n})+h[b_{1}(k_{1}(u_{n})-k_{1}(v_{n}))+b_{2}(k_{2}(u_{n})-k_{2}(v_{n}))+b_{3}(k_{3}(u_{n})-k_{3}(v_{n}))]+h\tau_{n}|$
$\leq|u_{n}-v_{n}|+h\left[b_{1}|k_{1}(u_{n})-k_{1}(v_{n})|+b_{2}|k_{2}(u_{n})-k_{2}(v_{n})|+b_{3}|k_{3}(u_{n})-k_{3}(v_{n})|\right]+h|\tau_{n}|$
@@
$\quad+\left.b_{3}hL\left(1+h\left(a_{31}+ha_{32}(1+hL)\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\right)\right)\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\right)\left|u_{n}-v_{n}\right|+h|\tau_{n}|$
$=\phi_{n}E_{n}+h|\tau_{n}|$
@@
$\phi_{n}$ $=1+b_{1}hL+b_{2}(1+hL)\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}+b_{3}hL\left(1+(hLa_{31}+hLa_{32}(1+hL))\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\right)$
$\leq 1+hL+b_{2}hL(1+hL)\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}+b_{3}hL\left(1+hLa_{31}+hLa_{32}(1+hL)\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\right)$
$\leq e^{hL}+b_{2}hLe^{hL}\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}+b_{3}hL{e^{hL}+hLe^{hL}}\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}$
$\leq e^{hL}\biggl{(}1+b_{2}hL\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\biggl{)}+b_{3}hL(1+hL)\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}$
$\leq e^{hL}(1+\mathcal{C}_{3}h)$
$\leq e^{h(L+\mathcal{C}_{3})}$
@@
$\mathcal{C}_{3}=\sup_{\begin{subarray}{c}0<h\leq b-a\\ n=0,\dots,N-1\end{subarray}}\frac{b_{2}L\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}+b_{3}L(1+hL)\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}{h},$
@@
$E_{n}\leq he^{(L+C_{2})T}\sum_{j=0}^{n-1}\|\tau\|_{\infty}=nhe^{(L+C_{2})(b-a)}\|\tau\|_{\infty}\leq(b-a)e^{(L+C_{2})(b-a)}\|\tau\|_{\infty}$
@@
$-f\left(t_{n},\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}\,\,hc_{2}k_{1}(v_{n})+\frac{v_{n}}{\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}}\right)\Biggl{|}$
@@
$-\frac{v_{n}}{\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}}\Biggl{|}$
$\leq L\biggl{|}\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}\,\,hc_{2}(f(u_{n})-f(v_{n}))+\frac{u_{n}-v_{n}}{\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}}\biggl{|}$
@@
$E_{n+1}$ $=|u_{n+1}-v_{n+1}|$
$=|(u_{n}-v_{n})+h[b_{1}(k_{1}(u_{n})-k_{1}(v_{n}))+b_{2}(k_{2}(u_{n})-k_{2}(v_{n}))]+h\tau_{n}|$
$\leq|u_{n}-v_{n}|+h[b_{1}|k_{1}(u_{n})-k_{1}(v_{n})|+b_{2}|k_{2}(u_{n})-k_{2}(v_{n})|]+h|\tau_{n}|$
$\leq|u_{n}-v_{n}|+hb_{1}L|u_{n}-v_{n}|+hb_{2}L\left|\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}\,\,hc_{2}L+\frac{1}{\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}}\right||u_{n}-v_{n}|+h|\tau_{n}|$
@@
$\phi_{n}$ $=1+b_{1}hL+hb_{2}L\left(\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\,\,Lhc_{2}+\frac{1}{\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}}\right)$
$\leq 1+b_{1}hL+hb_{2}L\left(\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\,\,Lhc_{2}+1\right)$
$=1+b_{1}hL+hb_{2}L\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\,\,Lha_{21}+hb_{2}L$
$=1+hb_{1}L+hb_{2}L\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}Lha_{21}+hL(1-b_{1})$
$=1+hL+hb_{2}L\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\,\,Lha_{21}$
@@
$\phi_{n}$ $\leq e^{hL}+\mathcal{C}_{2}he^{hL}$
$\leq(1+\mathcal{C}_{2}h)e^{hL}$
$\leq e^{h(L+\mathcal{C}_{2})}$

### v2

@@ 3.2 Three-stage fourth-order methods
$\tau_{n}=$ $\frac{u_{n+1}-u_{n}}{h}-(b_{1}k_{1}+b_{2}k_{2}+b_{3}k_{3}),$
$=$ $(1-b_{1}-b_{2}-b_{3})f+h\left[ff_{u}\left(\frac{1}{2}-a_{21}b_{2}-a_{31}b_{3}-a_{32}b_{3}\right)+f_{t}\left(\frac{1}{2}-b_{2}c_{2}-b_{3}c_{3}\right)+\frac{f^{2}}{2}\right]$
[…]
$\quad\left.\left.-a_{21}a_{32}b_{3}c_{3}\right)+ff_{utt}\left(\frac{1}{8}-\frac{a_{21}b_{2}c_{2}^{2}}{2}-\frac{a_{31}b_{3}c_{3}^{2}}{2}-\frac{a_{32}b_{3}c_{3}^{2}}{2}\right)+ff_{t}f_{uu}\left(\frac{1}{8}-a_{31}a_{32}b_{3}c_{2}-a_{32}^{2}b_{3}c_{2}\right)\right$.
$\quad\left.+f_{tt}f_{u}\left(\frac{1}{24}-\frac{a_{32}b_{3}c_{2}^{2}}{2}\right)+f_{t}f_{ut}\left(\frac{1}{8}-a_{32}b_{3}c_{2}c_{3}\right)+f_{ttt}\left(\frac{1}{24}-\frac{b_{2}c_{2}^{3}}{6}-\frac{b_{3}c_{3}^{3}}{6}\right)+\frac{a_{21}b_{2}c_{2}^{2}u_{n}\epsilon_{n2}^{2}ff_{uu}}{2}\right$.
$\quad\left.-\frac{a_{21}b_{2}c_{2}^{2}\epsilon_{n2}^{2}ff_{u}}{2}\right]+\mathcal{O}(h^{4}).$
@@
    $\begin{aligned} &a_{21}=c_{2}=\frac{1}{2},a_{31}=-1,a_{32}=2,c_{3}=1,b_{1}=\frac{1}{6},b_{2}=\frac{2}{3},b_{3}=\frac{1}{6},\\ &\epsilon^{2}_{n2}=\frac{-f^{2}f_{u}f_{uu}-ff_{t}f_{uu}+ff_{u}^{3}-ff_{u}f_{tu}+f_{t}f_{u}^{2}-f_{t}f_{tu}}{uff_{uu}-uf_{u}^{2}+uf_{tu}-ff_{u}},\\ &\epsilon^{2}_{n3}=-\epsilon^{2}_{n2}.\end{aligned}$
@@
[…]
@@
    $\begin{aligned} &a_{21}=c_{2}=1,a_{31}=\frac{1}{4},a_{32}=\frac{1}{4},c_{3}=\frac{1}{2},b_{1}=\frac{1}{6},b_{2}=\frac{1}{6},b_{3}=\frac{2}{3},\\ &\epsilon^{2}_{n2}=\frac{-ff_{t}f_{uu}-ff_{u}^{3}+ff_{u}f_{tu}-f_{t}f_{u}^{2}-f_{t}f_{tu}+f_{tt}f_{u}}{uff_{uu}+2vf_{u}^{2}+uf_{tu}-ff_{u}},\\ &\epsilon^{2}_{n3}=-\epsilon^{2}_{n2}.\end{aligned}$
@@
    $\begin{aligned} &a_{21}=c_{2}=\frac{1}{2},a_{31}=0,a_{32}=c_{3}=\frac{3}{4},c_{3}=1,b_{1}=\frac{2}{9},b_{2}=\frac{1}{3},b_{3}=\frac{4}{9},\\ &\epsilon^{2}_{n2}=\frac{\frac{1}{3}f^{3}f_{uuu}+f^{2}f_{tuu}+4ff_{u}^{3}+ff_{ttu}+4f_{t}f_{u}^{2}+\frac{1}{3}f_{ttt}}{uff_{uu}-4uf_{u}^{2}+uf_{tu}-ff_{u}},\\ &\epsilon^{2}_{n3}=-\frac{1}{3}\epsilon^{2}_{n2}.\end{aligned}$
4 Convergence of RBF Runge-Kutta Methods
Numerical methods derived in previous section can be applied on IVP if there exists a unique solution. IVP of the form (1.1) has a unique solution if $f(t,u)$ satisfies Lipschitz continuity with respect to second-coordinate that is $|f(t,u_{1})-f(t,u_{2})|\leq L|u_{1}-u_{2}|$, where $L$ is Lipschitz constant. Now, we prove the convergence of above proposed methods. Note that, we solve the IVP on an interval $[a,b]$ which is divided by uniform mesh length $h=\dfrac{b-a}{N}$ as $N+1$ points with mesh grid points $a=t_{0}<t_{1}<...<t_{N}=b$ where $t_{n}=a+nh,n=0,1,...,N.$ We also denote $u_{n}=u(t_{n})$ as exact solution and $v_{n}$ as numerical approximation for the proposed RBF Runge-Kutta schemes.
@@ 4 Convergence of RBF Runge-Kutta Methods
Suppose, $E_{n+1}$ be the error between $u_{n+1}$ and $v_{n+1}$ for fixed $t=t_{n}$. Now,
@@
$\leq L\biggl|\left(u_{n}+ha_{21}k_{1}(u_{n})\right)\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}-\left(v_{n}+ha_{21}k_{1}(v_{n})\right)\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}\biggl|,$
$\leq L\biggl|(u_{n}-v_{n})\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}+a_{21}hL(u_{n}-v_{n})\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\biggl|,$
@@
where $L$ is Lipschitz constant. Now, we compute error
$E_{n+1}$ $=|u_{n+1}-v_{n+1}|,$
$=|(u_{n}-v_{n})+h[b_{1}(k_{1}(u_{n})-k_{1}(v_{n}))+b_{2}(k_{2}(u_{n})-k_{2}(v_{n}))]+h\tau_{n}|,$
$\leq|u_{n}-v_{n}|+h[b_{1}|k_{1}(u_{n})-k_{1}(v_{n})|+b_{2}|k_{2}(u_{n})-k_{2}(v_{n})|]+h|\tau_{n}|,$
$\leq\left(1+hLb_{1}+b_{2}(1+hL)\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\right)|u_{n}-v_{n}|+h|\tau_{n}|,$
$=\phi_{n}E_{n}+h|\tau_{n}|,$
@@
$\phi_{n}$ $=1+b_{1}hL+b_{2}(1+a_{21}hL)\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}},$
$\leq 1+hL+e^{hL}\mathcal{C}_{2}h,$
$\leq e^{hL}+e^{hL}\mathcal{C}_{2}h,$
$=e^{hL}(1+\mathcal{C}_{2}h),$
$=e^{h(L+\mathcal{C}_{2})},$
@@
$\mathcal{C}_{2}=\sup_{\begin{subarray}{c}0<h\leq b-a\\ n=0,\dots,N-1\end{subarray}}\frac{\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}}{h}.$
@@
$\qquad-f\left(t_{n},(v_{n}+h(a_{31}k_{1}(v_{n})+a_{32}k_{2}(v_{n}))\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\right)\bigg|,$
$\leq L\left|(u_{n}-v_{n})+\left(ha_{31}L|u_{n}-v_{n}|+ha_{32}L(1+hL)\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}|u_{n}-v_{n}|\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\right)\right|,$
@@
where $L$ is Lipschitz constant.
$E_{n+1}$ $=|u_{n+1}-v_{n+1}|,$
$=|(u_{n}-v_{n})+h[b_{1}(k_{1}(u_{n})-k_{1}(v_{n}))+b_{2}(k_{2}(u_{n})-k_{2}(v_{n}))+b_{3}(k_{3}(u_{n})-k_{3}(v_{n}))]+h\tau_{n}|,$
$\leq|u_{n}-v_{n}|+h\left[b_{1}|k_{1}(u_{n})-k_{1}(v_{n})|+b_{2}|k_{2}(u_{n})-k_{2}(v_{n})|+b_{3}|k_{3}(u_{n})-k_{3}(v_{n})|\right]+h|\tau_{n}|,$
@@
$\quad+\left.b_{3}hL\left(1+h\left(a_{31}+ha_{32}(1+hL)\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\right)\right)\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\right)\left|u_{n}-v_{n}\right|+h|\tau_{n}|,$
$=\phi_{n}E_{n}+h|\tau_{n}|.$
@@
$\phi_{n}$ $=1+b_{1}hL+b_{2}(1+hL)\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}+b_{3}hL\left(1+(hLa_{31}+hLa_{32}(1+hL))\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\right),$
$\leq 1+hL+b_{2}hL(1+hL)\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}+b_{3}hL\left(1+hLa_{31}+hLa_{32}(1+hL)\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}\right),$
$\leq e^{hL}+b_{2}hLe^{hL}\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}+b_{3}hL{e^{hL}+hLe^{hL}}\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}},$
$\leq e^{hL}\biggl(1+b_{2}hL\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\biggl)+b_{3}hL(1+hL)\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}},$
$\leq e^{hL}(1+\mathcal{C}_{3}h),$
$\leq e^{h(L+\mathcal{C}_{3})},$
@@
$\mathcal{C}_{3}=\sup_{\begin{subarray}{c}0<h\leq b-a\\ n=0,\dots,N-1\end{subarray}}\frac{b_{2}L\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}+b_{3}L(1+hL)\sqrt{1+\epsilon_{n3}^{2}c_{3}^{2}h^{2}}}{h}.$
@@
$E_{n}\leq he^{(L+C_{2})T}\sum_{j=0}^{n-1}\|\tau\|_{\infty}=nhe^{(L+C_{2})(b-a)}\|\tau\|_{\infty}\leq(b-a)e^{(L+C_{2})(b-a)}\|\tau\|_{\infty},$
@@
$\qquad-f\left(t_{n},\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}\,\,hc_{2}k_{1}(v_{n})+\frac{v_{n}}{\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}}\right)\Biggl|,$
@@
$-\frac{v_{n}}{\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}}\Biggl|,$
$\leq L\biggl|\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}\,\,hc_{2}(f(u_{n})-f(v_{n}))+\frac{u_{n}-v_{n}}{\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}}\biggl|,$
@@
$E_{n+1}$ $=|u_{n+1}-v_{n+1}|,$
$=|(u_{n}-v_{n})+h[b_{1}(k_{1}(u_{n})-k_{1}(v_{n}))+b_{2}(k_{2}(u_{n})-k_{2}(v_{n}))]+h\tau_{n}|,$
$\leq|u_{n}-v_{n}|+h[b_{1}|k_{1}(u_{n})-k_{1}(v_{n})|+b_{2}|k_{2}(u_{n})-k_{2}(v_{n})|]+h|\tau_{n}|,$
$\leq|u_{n}-v_{n}|+hb_{1}L|u_{n}-v_{n}|+hb_{2}L\left|\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}\,\,hc_{2}L+\frac{1}{\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}}\right||u_{n}-v_{n}|+h|\tau_{n}|,$
@@
$\phi_{n}$ $=1+b_{1}hL+hb_{2}L\left(\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\,\,Lhc_{2}+\frac{1}{\sqrt{(1+\epsilon_{n2}^{2}c_{2}^{2}h^{2})}}\right),$
$\leq 1+b_{1}hL+hb_{2}L\left(\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\,\,Lhc_{2}+1\right),$
$=1+b_{1}hL+hb_{2}L\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\,\,Lha_{21}+hb_{2}L,$
$=1+hb_{1}L+hb_{2}L\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}Lha_{21}+hL(1-b_{1}),$
$=1+hL+hb_{2}L\sqrt{1+\epsilon_{n2}^{2}c_{2}^{2}h^{2}}\,\,Lha_{21},$
@@
$\phi_{n}$ $\leq e^{hL}+\mathcal{C}_{2}he^{hL},$
$\leq(1+\mathcal{C}_{2}h)e^{hL},$
$\leq e^{h(L+\mathcal{C}_{2})}.$